OpenStudy (anonymous):

What adds to be 3 but multiplies to be -18

OpenStudy (anonymous):

6 & -3

OpenStudy (anonymous):

6 and -3. 6+(-3)=3 6*(-3)=-18

OpenStudy (anonymous):

what is the answer to 5x^2+24x-36=0